OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation. y = 6 when x = 8

OpenStudy (anonymous):

y = k/x 6 = k/8 6*8 = k 48 = k k = 48

OpenStudy (anonymous):

48/x?

OpenStudy (anonymous):

or x/48

OpenStudy (anonymous):

the equation is y= 48/x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np